# Efficient optimization of neural network backflow for ab-initio quantum chemistry

### v1

@@ Appendix B Learned Orbital Envelope Parameters
This section illustrates what the orbital envelope [Eq. (II.5.3)] learns during training by visualizing the distribution of its trainable parameters for the last four molecules listed in Table 1. All $\alpha$ values are initially set to a small value (0.01) to prevent biasing the ansatz. As shown in Figure 6, after training, the $\alpha$ values associated with inner electrons become noticeably larger than those for outer electrons, reflecting the intuitive physical expectation that inner electrons remain closer to the nucleus in lower-energy orbitals.
Figure 6: Distribution of the orbital envelope parameters $\alpha_{i}$ after training for CH4O, Li2O, C2H4O, and C2H4O2 in the STO-3G basis with CCSD natural orbitals. The y-axis shows the absolute value of $\alpha_{i}$, as only its magnitude affects the envelope per Eq. (II.5.3).

### v2

@@ Appendix B Learned Orbital Envelope Parameters
This section illustrates what the orbital envelope [Eq. (16)] learns during training by visualizing the distribution of its trainable parameters for the last four molecules listed in Table 1. All $\alpha$ values are initially set to a small value (0.01) to prevent biasing the ansatz. As shown in Figure 8, after training, the $\alpha$ values associated with inner electrons become noticeably larger than those for outer electrons, reflecting the intuitive physical expectation that inner electrons remain closer to the nucleus in lower-energy orbitals.
Figure 8: Distribution of the orbital envelope parameters $\alpha_{i}$ after training for CH4O, Li2O, C2H4O, and C2H4O2 in the STO-3G basis with CCSD natural orbitals. The y-axis shows the absolute value of $\alpha_{i}$, as only its magnitude affects the envelope per Eq. (16).